# Emotional Support with LLM-based Empathetic Dialogue Generation

### v1

@@ 1 Introduction
Emotional Support Conversation (ESC) aims to alleviate individuals’ emotional distress and provide effective emotional support through conversational interactions[17, 18, 21]. With the increasing prevalence of mental health issues globally, ESC has demonstrated significant potential as an intervention tool in various domains, including mental health support, customer service, and interpersonal interactions[7, 8, 9, 14, 19, 20, 24, 29]. ESC systems not only need to process information from everyday conversations but must also focus on understanding and addressing users’ emotional needs, providing targeted emotional responses. As such, the design of emotional support dialogue systems faces several challenges, such as accurately understanding subtle emotional fluctuations in users, generating emotionally authentic and constructive responses, and designing support strategies that are adaptable to different contexts and user needs[33].
Despite significant advancements in sentiment analysis and emotional response generation technologies in recent years, existing emotional support dialogue systems still have considerable limitations. Many systems struggle to effectively capture the nuances and complexity of emotions, and the generated responses often lack true emotional resonance and practical help[11, 15]. This represents a critical area for continued exploration and optimization in the ESC field. To further advance research in this domain, the NLPCC 2025 Task 8 has organized an Emotional Support Conversation evaluation task, aimed at assessing different models’ capabilities to understand and respond to users’ emotional needs, providing a unified dataset and evaluation standards.
In this evaluation task, our team adopted a fine-tuning approach based on large-scale language models (LLMs) and achieved second place [13, 25]. LLMs were selected as the core models due to their remarkable capabilities in natural language understanding and generation, making them ideal for the ESC task. By fine-tuning these models on a specialized emotional support dialogue dataset, we aimed to better align the models with the specific demands of ESC tasks, enabling them to generate empathetic and contextually appropriate responses that cater to users’ emotional needs. Specifically, we employed the qwen2.5-72B-instruct and qwen2.5-7B-Instruct models, applying both low-rank adaptation (LoRA) fine-tuning and full-parameter fine-tuning strategies to further improve performance. Our experimental results demonstrate that appropriate fine-tuning techniques can significantly enhance the performance of models in emotional support conversation tasks.
@@ 2.1 Rule-Based Approaches
Rule-based ESC systems typically rely on manually designed rules and templates to generate supportive responses. These systems identify keywords or patterns in the user’s input and select appropriate responses from a predefined list. For example, a system may recognize a specific topic mentioned by the user and respond using a predefined template or scripted message [4]. Early conversational agents, such as ELIZA, were built using rule-based methods, providing simple emotional support through scripted dialogue. However, rule-based approaches have limitations in capturing the subtle nuances of human emotions, resulting in responses that may feel rigid and predictable [1, 27]. While these methods offer certain advantages in terms of safety and controllability—ensuring that inappropriate responses are avoided—they lack the flexibility and depth required to provide truly empathetic support. The inability to deeply understand the user’s emotional state remains a major constraint.
@@ 2.2 Retrieval-Based Approaches
Retrieval-based ESC systems generate responses by selecting the most relevant reply from a predefined corpus. These methods often rely on keyword matching or machine learning techniques to find the most similar response to the user’s input. Retrieval-based systems have advantages in ensuring the coherence and computational efficiency of responses [6, 16, 23, 26]. Some studies have also explored incorporating techniques such as emotional clustering and explainable AI into retrieval-based ESC systems to enhance their effectiveness. However, the performance of retrieval-based methods is heavily dependent on the quality and coverage of the predefined corpus [30, 32]. As a result, they may struggle to provide appropriate responses to novel situations or scenarios that were not anticipated in the corpus.
@@ 2.3 Generation-Based Approaches
Generation-based ESC systems, particularly those employing neural networks and large-scale language models (LLMs), are capable of generating entirely new responses. These approaches train models on vast amounts of conversational data, enabling them to understand context and generate natural language responses [2, 3, 12, 28]. However, the use of generative methods in ESC tasks introduces several challenges, such as maintaining conversational coherence and ensuring that the generated emotional responses are appropriate [10, 22]. With the rapid development of LLMs in recent years, there has been increasing exploration into leveraging LLMs and fine-tuning techniques to build more effective ESC systems. LLMs, with their strong generalization ability and capacity to learn complex patterns from large datasets, hold significant promise for enhancing emotional support dialogue systems.
In addition to the aforementioned categories, many studies have focused on how to integrate emotional intelligence and empathy into dialogue systems. This includes research on modeling the user’s emotional state and generating responses that demonstrate empathy [5, 31]. Some studies have emphasized the importance of considering the help-seeker’s personal traits when providing effective support. Furthermore, generating inappropriate responses remains a critical issue, and researchers are actively exploring various techniques to mitigate this problem. Building genuinely empathetic ESC systems requires not only understanding the user’s emotions but also accounting for their specific circumstances, personality, and the potential for inappropriate responses.

### v2

@@ 1 Introduction
Emotional Support Conversation (ESC) aims to alleviate individuals’ emotional distress and provide effective emotional support through conversational interactions[song2015detecting, song2019concept, song2021emotion]. With the increasing prevalence of mental health issues globally, ESC has demonstrated significant potential as an intervention tool in various domains, including mental health support, customer service, and interpersonal interactions[jiang2025utterance, jiang2025label, kang2024can, liu2021towards, shi2025muddit, liu2023icsberts, song2021emotional, song2020sentiment, wang2024towards, ouyang2023mining, xiong2024dual, yao2020session, song2015classifying, wu2025table, zhang2024lemur, song2020tcnn, song2020two, wu2025mr, li2024scalable, ning2023ump, zhao2023muse, song2015recommending]. ESC systems not only need to process information from everyday conversations but must also focus on understanding and addressing users’ emotional needs, providing targeted emotional responses. As such, the design of emotional support dialogue systems faces several challenges, such as accurately understanding subtle emotional fluctuations in users, generating emotionally authentic and constructive responses, and designing support strategies that are adaptable to different contexts and user needs[zhou2023facilitating].
Despite significant advancements in sentiment analysis and emotional response generation technologies in recent years, existing emotional support dialogue systems still have considerable limitations. Many systems struggle to effectively capture the nuances and complexity of emotions, and the generated responses often lack true emotional resonance and practical help[li2024helpful, shao2025ai]. This represents a critical area for continued exploration and optimization in the ESC field. To further advance research in this domain, the NLPCC 2025 Task 8 has organized an Emotional Support Conversation evaluation task, aimed at assessing different models’ capabilities to understand and respond to users’ emotional needs, providing a unified dataset and evaluation standards.
In this evaluation task, our team adopted a fine-tuning approach based on large-scale language models (LLMs) and achieved second place [liu2024icsplms, wang2024telechat]. LLMs were selected as the core models due to their remarkable capabilities in natural language understanding and generation, making them ideal for the ESC task. By fine-tuning these models on a specialized emotional support dialogue dataset, we aimed to better align the models with the specific demands of ESC tasks, enabling them to generate empathetic and contextually appropriate responses that cater to users’ emotional needs. Specifically, we employed the qwen2.5-72B-instruct and qwen2.5-7B-Instruct models, applying both low-rank adaptation (LoRA) fine-tuning and full-parameter fine-tuning strategies to further improve performance. Our experimental results demonstrate that appropriate fine-tuning techniques can significantly enhance the performance of models in emotional support conversation tasks.
@@ 2.1 Rule-Based Approaches
Rule-based ESC systems typically rely on manually designed rules and templates to generate supportive responses. These systems identify keywords or patterns in the user’s input and select appropriate responses from a predefined list. For example, a system may recognize a specific topic mentioned by the user and respond using a predefined template or scripted message [chu2024towards]. Early conversational agents, such as ELIZA, were built using rule-based methods, providing simple emotional support through scripted dialogue. However, rule-based approaches have limitations in capturing the subtle nuances of human emotions, resulting in responses that may feel rigid and predictable [alazraki2021empathetic, yu2024experimental]. While these methods offer certain advantages in terms of safety and controllability—ensuring that inappropriate responses are avoided—they lack the flexibility and depth required to provide truly empathetic support. The inability to deeply understand the user’s emotional state remains a major constraint.
@@ 2.2 Retrieval-Based Approaches
Retrieval-based ESC systems generate responses by selecting the most relevant reply from a predefined corpus. These methods often rely on keyword matching or machine learning techniques to find the most similar response to the user’s input. Retrieval-based systems have advantages in ensuring the coherence and computational efficiency of responses [jia2023knowledge, song2017intention, wang2022topicks, xu2024dynamic]. Some studies have also explored incorporating techniques such as emotional clustering and explainable AI into retrieval-based ESC systems to enhance their effectiveness. However, the performance of retrieval-based methods is heavily dependent on the quality and coverage of the predefined corpus [zhao2023transesc, zheng2024self]. As a result, they may struggle to provide appropriate responses to novel situations or scenarios that were not anticipated in the corpus.
@@ 2.3 Generation-Based Approaches
Generation-based ESC systems, particularly those employing neural networks and large-scale language models (LLMs), are capable of generating entirely new responses. These approaches train models on vast amounts of conversational data, enabling them to understand context and generate natural language responses [chen2024cauesc, cheng2023pal, li2024tele, zhao2024autograph]. However, the use of generative methods in ESC tasks introduces several challenges, such as maintaining conversational coherence and ensuring that the generated emotional responses are appropriate [kluwer2011chatbots, vanel2023survey]. With the rapid development of LLMs in recent years, there has been increasing exploration into leveraging LLMs and fine-tuning techniques to build more effective ESC systems. LLMs, with their strong generalization ability and capacity to learn complex patterns from large datasets, hold significant promise for enhancing emotional support dialogue systems.
In addition to the aforementioned categories, many studies have focused on how to integrate emotional intelligence and empathy into dialogue systems. This includes research on modeling the user’s emotional state and generating responses that demonstrate empathy [he2024telechat, zhao2025fisminess]. Some studies have emphasized the importance of considering the help-seeker’s personal traits when providing effective support. Furthermore, generating inappropriate responses remains a critical issue, and researchers are actively exploring various techniques to mitigate this problem. Building genuinely empathetic ESC systems requires not only understanding the user’s emotions but also accounting for their specific circumstances, personality, and the potential for inappropriate responses.